OpenStudy (precal):

Does the Mean Value Theorem apply to f(x)=(x+1)/x from [1,3]

OpenStudy (precal):



OpenStudy (amistre64):

is it continuous from 1 to 3?

OpenStudy (precal):

yes but continuity does not imply differentiablility

OpenStudy (amistre64):

but it helps ....

OpenStudy (precal):

I know..

OpenStudy (precal):

I took the derivative and got \[f \prime(x)=\frac{-1}{x^2}\]

OpenStudy (dumbcow):

nevermind

OpenStudy (amistre64):

you have smooth curves, well behaved along intervals except for in that one spot ...

OpenStudy (precal):

I am looking for the c value that the Mean Value Thm guarantees

OpenStudy (amistre64):

what your slope from 1 to 3?

OpenStudy (precal):

-1/3

OpenStudy (amistre64):

the -x^(-2) Ds to get 2x^(-3)

OpenStudy (amistre64):

when 2x^(-3) = -1/3 you got it

OpenStudy (precal):

I set that equal to the derivative to solve for x

OpenStudy (amistre64):

when x^3 = 6 it looks like

OpenStudy (amistre64):

-6 as the cace may be

OpenStudy (amistre64):

is the cbrt (-6) in the interval?

OpenStudy (precal):

I got \[x=\sqrt{3}\] for x. Did I do something wrong?

OpenStudy (amistre64):

i mighta did something wrong with these fat little fingers ...

OpenStudy (dumbcow):

i also get sqrt(3) :)

OpenStudy (amistre64):

your initial function is -x^(-2) right? 2x^(-3) is the derivative ... right?

OpenStudy (amistre64):

i see what im doing, im confusing a few posts in this thing ...

OpenStudy (amistre64):

ill go with sqrt(3) by default then :)

OpenStudy (precal):

ok thanks, I think I will stick with the math and forget the math movies.

OpenStudy (amistre64):

everyone needs a hobby :)

OpenStudy (precal):

yes, I just don't have time to read anything other than math books. Just not enough time.......